### v4

# AccidentSim: Generating Vehicle Collision Videos with Physically Realistic Collision Trajectories from Real-World Accident Reports

Abstract
Collecting real-world vehicle accident videos for autonomous driving research is challenging due to their rarity and complexity. While existing video generation methods can produce visually realistic content, they often lack physical realism and fail to generate accurate post-collision vehicle trajectories. In this paper, we introduce AccidentSim, a framework designed to generate physically realistic vehicle collision videos. Specifically, AccidentSim extracts and utilizes physical cues and contextual information from real-world accident reports. It then uses a reliable physical simulator to replicate post-collision vehicle trajectories based on the information from these reports, creating a collision trajectory dataset. This dataset is subsequently used to fine-tune a language model, AccidentLLM, enabling it to predict novel, physically consistent post-collision trajectories across various driving scenarios based on user-provided accident descriptions. Finally, we synthesize collision videos by compositing the predicted trajectories onto vehicles with rendered backgrounds. Experimental results demonstrate that videos generated by AccidentSim excel in both visual and physical realism, and incorporating these videos into autonomous system training can effectively reduce collision rates.
Figure 1: AccidentSim generates vehicle collision videos from user-provided accident descriptions, producing physically realistic collision trajectories of vehicles across diverse collision and road types, all according to the user-specified viewpoints. Particularly, AccidentSim extracts physical cues and contextual information from real-world accident reports and feeds them into a physical simulator to generate physically constrained collision trajectories. These trajectories are then used to fine-tune a language model, termed AccidentLLM, enabling it to predict novel, physically consistent trajectories directly from the collision information extracted from the user-provided accident descriptions, extending beyond the collision scenarios contained in the original accident reports.
I Introduction
Research and development in autonomous driving critically depend on access to diverse and comprehensive training data to ensure robust performance across various driving scenarios. However, existing datasets, such as KITTI [17], Cityscapes [8], nuScenes [4], Waymo [44], and BDD100K [61], primarily capture normal driving conditions, lacking safety-critical events such as vehicle collisions. This data imbalance arises from the inherent difficulty and prohibitive cost of collecting rare but safety-critical scenarios in real-world environments. This phenomenon is known as the Curse of Rarity (CoR) [29], which refers to the low occurrence rate of safety-critical events in high-dimensional spaces. In autonomous driving systems, CoR significantly limits the ability of deep learning models to learn effective responses to these rare but crucial situations, ultimately undermining the overall reliability and safety of such systems.
With recent advances in video generation models, synthesizing vehicle collision videos has become a feasible, low-cost, and promising alternative to expensive real-world data collection. Such generated videos, if physically realistic, could augment existing datasets and enhance the training of autonomous driving systems, ultimately improving reliability and safety in critical scenarios. However, current video generation methods, including state-of-the-art models such as Kling [27], Gen-4 [40], and Jimeng [3], as well as specialized driving-scenario generators like DriveDreamer [49], Panacea [51], Cosmos [1], and HoloDrive [53], still struggle to produce physically realistic vehicle collisions. In particular, these approaches often fail to reproduce physically realistic collision trajectories that conform to real-world physical constraints and dynamics. This limitation arises because these video generation methods are not explicitly designed to incorporate essential physical cues and constraints. For instance, existing models typically overlook key physical factors, such as vehicle momentum, impulse, and mass distribution, that are critical for accurately modeling real-world collision behaviors. Consequently, while the videos generated by these models may be visually realistic, they lack physical realism, failing to capture the true underlying dynamics of vehicle collisions.
To address these challenges, we propose AccidentSim, a framework that introduces a novel approach to generating physically realistic vehicle collision videos based on real-world accident reports. For instance, accident reports from the National Motor Vehicle Crash Causation Survey (NMVCCS) [34] in the United States document actual vehicle collisions across diverse vehicle models, collision types, road types, and traffic conditions. AccidentSim leverages the rich physical cues and contextual information embedded in such reports to synthesize vehicle collision videos with physically realistic post-collision trajectories. As illustrated in Figure 1, AccidentSim first extracts key physical cues, such as vehicle speed, collision angle, and mass, together with contextual information, including road type and collision type, from the accident reports. These cues serve as the physical foundation for constructing realistic collision simulations. The extracted physical cues are then fed into a physical simulator, such as CARLA [10], which enforces corresponding physical constraints to ensure that the simulated post-collision vehicle trajectories adhere to real-world dynamics. However, physical simulators have inherent limitations. For example, physical simulators are typically computationally expensive and require fully specified physical parameters, making them unable to directly respond to user-provided accident descriptions for generating diverse post-collision trajectories. To overcome these challenges, we first construct a dataset from the collision trajectories generated by the simulator using real-world accident reports. This dataset is then used to fine-tune a large language model, termed AccidentLLM, within the AccidentSim framework. As a result, AccidentLLM can generate novel, physically realistic collision trajectories directly from the collision information extracted from the user-provided accident descriptions, extending beyond the scenarios captured in existing accident reports. In this way, AccidentLLM effectively bypasses the expensive simulation process through physics distillation, learning to translate under-specified semantic inputs into physically realistic trajectories efficiently and realistically. Finally, AccidentSim synthesizes collision videos by compositing the predicted trajectories onto vehicles with rendered backgrounds, according to user-specified viewpoints.
In summary, AccidentSim integrates physically grounded cues with physical constraints to generate videos that faithfully replicate diverse real-world vehicle collisions. By bridging the gap between synthetic and real-world scenarios, AccidentSim offers a scalable solution to the scarcity of physically realistic collision data for training and evaluating autonomous driving systems.
Our contributions are three-fold:
  • We propose AccidentSim, a framework that extracts and leverages physical cues and contextual information from real-world accident reports, to generate vehicle collision videos with physical simulation, ensuring physically realistic post-collision trajectories.
  • We develop AccidentLLM, a language-driven module fine-tuned on simulated collision trajectories for pre-collision planning and post-collision prediction, enabling the generation of novel, physically realistic trajectories directly from user-provided accident descriptions.
  • We demonstrate AccidentSim’s generalization across diverse road types, collision types, and driving scenarios, validating its adaptability and applicability to a wide range of real-world conditions.
II Related Work
II-A Safety-Critical Scenarios and Driving Datasets
The foundation of autonomous driving research relies on large-scale datasets, such as KITTI [17], Cityscapes [8], nuScenes [4], Waymo Open Dataset [44], and BDD100K [61]. While providing massive normal driving trajectories, they inherently lack safety-critical vehicle collisions. To address this, early efforts constructed specialized accident datasets, such as CADP [42], ISSAFE [63], and CCD [2], which are enriched with spatio-temporal and environmental annotations. Recently, advanced paradigms like EQ-TAA [12] employ video diffusion to synthesize pseudo-accident pairs to actively mitigate data bias. However, these datasets and frameworks are bottlenecked by fixed viewpoints and prohibitive acquisition costs. More importantly, while excelling at observational anticipation, they fundamentally lack the capability to dynamically synthesize novel crash scenarios. To bypass real-world data collection, adversarial scenario generation and trajectory optimization have been widely explored. Methods like AdvSim [47], Learning-to-Collide [9], and CAT [65] employ adversarial training for challenging pre-collision scenarios, while STRIVE [38] and others [66] optimize latent traffic models evaluated in platforms like SafeBench [55]. However, these approaches focus solely on pre-collision obstacle avoidance. They fail to simulate actual physical impacts and post-collision dynamics, leaving a critical gap in fully replicating the accident lifecycle.
II-B Trajectory Prediction and Language-Guided Generation
Accurate vehicle motion forecasting is core to driving simulation. Traditional multi-agent trajectory prediction models [43, 25, 35, 13] rely heavily on historical states, latent sequence modeling [18], and high-definition maps. To explicitly capture temporal dependencies across multiple trajectories, recent efforts integrate spatio-temporal sequence fusion to comprehensively model social interactions among agents [48]. Recently, language-guided traffic generation driven by large language models (LLMs) [11] has emerged. Frameworks like LCTGen [45], CTG++ [68], TrafficGen [14], and Traj-LLM [58] control traffic simulation via natural language, while MotionLM [41] and MotionCLIP [46] frame forecasting directly as a language modeling task. To bridge high-level reasoning and low-level control, recent studies distill multi-modal LLM knowledge into efficient vision-based planners [20], enhancing generalizability in critical scenarios without prohibitive inference costs. Alternatively, to manage online querying overhead, AdaDrive [67] proposes an adaptive collaborative system. By dynamically invoking the LLM exclusively during complex scenarios, it balances high-level reasoning with fast conventional planners.
II-C Physical Simulation and Collision Modeling
Early analytical models [37, 69] utilized momentum, restitution, and friction to predict post-impact motion. Modern research integrates these principles into high-fidelity 3D simulators like CARLA [10], LGSVL [39], and UniSim [59] to enforce strict physical constraints. While these engines provide mathematically rigorous forward kinematics, integrating them into data-driven generative pipelines remains challenging. Recently, SoVAR [19] and AC3R [24] parsed real-world police reports from NMVCCS [34] to reconstruct accidents in simulators. Concurrently, generative architectures—such as combining Variational Autoencoders with Wasserstein GANs (VAE-WGAN)—have been integrated into end-to-end Deep Reinforcement Learning (DRL) frameworks to efficiently encode video features and explicitly minimize vehicle collision rates in simulators like CARLA [26]. However, pure physical simulations are computationally expensive, require meticulous parameter tuning, and cannot dynamically generate novel scenarios from under-specified prompts. AccidentSim resolves this by using simulators solely to construct a high-quality dataset. Distilling this data into AccidentLLM enables real-time, physics-aligned trajectory generation without the overhead of online physical rendering.
II-D Controllable Video Generation and World Models
Diffusion models have profoundly transformed driving simulation. Foundation models (e.g., Kling 2.0 [27], Gen-4 [40], Jimeng 3.0 [3], CogVideoX [60]) establish new photorealism benchmarks. Tailored for autonomous driving, generative world models [49, 28, 16, 51, 53, 33, 1] synthesize videos conditioned on layouts, text, or 3D geometry. Concurrently, NeRF and 3DGS advancements [54, 64, 31, 50, 70, 57] enable editable large-scale dynamic scene simulation. To further enhance 3D perception and dataset diversity, hardness-aware scene synthesis [62] adaptively composes 3D objects and backgrounds.
Moving beyond normal conditions, frameworks like Challenger [56] utilize physics-aware trajectory refinement to synthesize aggressive adversarial scenarios (e.g., cut-ins). To mitigate computational overhead in high-resolution multi-view synthesis, DriveScape [52] efficiently fuses multi-view features with 3D road structures. Moreover, self-consistent generative networks [30] rigidly preserve spatial consistency across diverse viewpoints with large camera baselines, bypassing explicit geometric rectification. Finally, MagicDrive-V2 [15] achieves high-resolution, long video generation with precise geometric control, while recent disentangled fine-tuning strategies [6] prevent attribute confusion by accurately binding specific motions and visual concepts to individual interacting agents in complex multi-vehicle scenarios.
Figure 2: Architecture of AccidentSim: AccidentSim extracts physical cues, such as vehicle speeds and collision types, along with contextual information like road types, from accident reports. This extracted information is then fed into a physical simulator, such as CARLA, to generate collision trajectories, which form a dataset. The dataset is subsequently used to fine-tune a language model, AccidentLLM (based on the Llama model), enabling it to adapt to various scenarios from user-provided accident descriptions and generate novel trajectories. Building on this foundation, AccidentSim leverages user accident descriptions to plan pre-collision trajectories, generating both the pre-collision trajectory and relevant collision information, such as vehicle speeds, collision angles, and more. The collision information is then processed by AccidentLLM, which produces novel, physically realistic post-collision trajectories, resulting in complete collision trajectories. These realistic trajectories are composited onto vehicles to generate the foreground, which is then combined with rendered backgrounds to produce collision videos across various scenarios.
III METHODOLOGY
AccidentSim first utilizes a large language model to extract key vehicle motion information (such as vehicle models, collision types, and intended video viewpoint movements) from the user-provided accident description. This extracted information is then processed by two core modules: the Pre-Collision Trajectory Planning module, which generates the trajectories leading to the collision, and the Post-Collision Prediction module, which generates the trajectories after the collision. Together, these modules produce a complete vehicle collision trajectory. Subsequently, in the following Collision Video Generation phase, the Foreground Processing module integrates the generated collision trajectory with user-specified vehicle models to produce foreground images. In the Background Processing module, background images are synthesized based on the user-specified viewpoint movement. Finally, the Video Composition module merges the foreground and background images to generate a vehicle collision video.
III-A Pre-Collision Trajectory Planning
At this stage, the proposed algorithm of AccidentSim produces pre-collision trajectories leading up to the collision, along with associated information such as vehicle speed. It operates across three primary phases. First, in the Initial State Extraction phase, essential vehicle and collision parameters are extracted from the user-provided accident description to establish the initial state for each vehicle and define the collision target. Second, in the Lane and Path Selection phase, the algorithm identifies suitable lanes based on vehicle dynamics and the scene map, forming potential lane combinations to ensure feasible paths for collision trajectory planning. Third, in the Trajectory Generation and Validation phase, valid pre-collision trajectories are generated using the selected lane combinations, which are then evaluated against collision criteria to yield a final set of feasible pre-collision trajectories.
III-A1 Initial State Extraction
In this phase, the proposed algorithm initiates by extracting essential vehicle and collision parameters from the vehicle motion information $\mathcal{I}$, which includes the dynamics of all vehicles involved in the scenario. Using this data, the algorithm isolates variables such as the collision point $\mathcal{C}$, vehicle count $n$, distances $D$, velocities $V$, accelerations $A$, and orientation angles $\theta$. This extraction process establishes a comprehensive view of each vehicle’s current state and potential collision zones. The algorithm then determines the starting points $P_{0}$ for each vehicle by factoring in directional and dynamic properties.
III-A2 Lane and Path Selection
This phase focuses on selecting suitable lanes within the scene map $\mathcal{M}$ to enable feasible trajectory planning. The proposed algorithm begins by parsing the scene map to identify available lanes $L$, ensuring that each candidate lane is assessed for compatibility with the vehicle’s current orientation $\theta$ and collision constraints $\mathcal{C}$. A subset of valid lanes $\mathcal{L}$ is then assembled, containing only those that meet the criteria defined by vehicle orientation and other relevant dynamics. After identifying valid lanes, the algorithm generates various lane combinations $\mathcal{G}$ by analyzing lane connectivity and spatial feasibility in relation to the parameters $n$, $D$, $V$, AA, and $\theta$.
III-A3 Trajectory Generation and Validation
This phase involves generating and validating trajectories based on the preselected lane combinations. For each lane combination $G_{i}$ in $\mathcal{G}$, the proposed algorithm checks whether it meets specific collision criteria using the collision parameters $\mathcal{C}$. Upon identifying a valid lane combination $G_{s}$ that satisfies these conditions, it proceeds to generate the trajectory paths $\mathcal{T}$. This process incorporates the starting points $P_{0}$ and valid lane combination $G_{s}$ to form trajectories that are both dynamically feasible and collision-aware. Each trajectory is then evaluated to ensure it conforms to the criteria established for pre-collision scenarios. The final output consists of the validated set of trajectories $\mathcal{T}$ and the associated imminent collision information $\mathcal{S}$. Subsequently, the imminent collision information $\mathcal{S}$ is provided as input to our “AccidentLLM”, which predicts novel post-collision trajectories (details provided in the Post-Collision Trajectory Prediction section).
III-B Post-Collision Trajectory Prediction
To address the challenge of collecting real-world post-collision trajectories, we leverage physical cues extracted from accident reports to train a customized version of the Llama model [11]. The resulting model, termed AccidentLLM, is capable of responding to the collision information extracted from the user-provided accident descriptions, and predicting novel post-collision trajectories that are adaptable to various road conditions.
As shown in Figure 2 (c), the proposed prediction is structured into three stages: (1) Accident Report Extraction, where collision-specific information is extracted from accident reports; (2) Collision Scenario Simulation, which generates physics-constrained post-collision trajectories; and (3) Fine-tuning AccidentLLM, where the extracted attributes and simulated scenarios are used to tailor the model for accurate trajectory prediction.
III-B1 Accident Report Extraction
The real-world accident reports used in our work, sourced from NMVCCS [34], provide critical data for supporting collision simulations, encompassing pre-accident environmental conditions, collision dynamics, and post-accident outcomes. Prior work, such as SoVAR [19], employs a template-based approach combined with language models to extract key information from accident reports. However, due to the inherent variability and complexity of natural language descriptions in these reports, a single-template strategy lacks the flexibility to adapt to diverse accident scenarios. To overcome this limitation, we employ a Llama model [11] with a few-shot learning strategy to automatically extract essential information from accident reports, including pre-accident environmental conditions, road features, and driver behaviors; in-collision details such as vehicle positioning, speed, and collision type; and post-accident information including emergency responses.
III-B2 Collision Scenario Simulation
In the accident simulation phase, we developed a comprehensive scenario reconstruction pipeline to accurately recreate real-world collision scenarios and construct a high-quality collision dataset for the subsequent fine-tuning of AccidentLLM. The process begins by partitioning the extracted accident report information into vehicle-specific information, such as velocity and location, and road-related details. Using RoadRunner [32], we construct high-fidelity 3D road models, which are seamlessly integrated into the CARLA simulation environment. CARLA’s advanced vehicle navigation capabilities, combined with the extracted vehicle dynamics and physical parameters such as friction and gravity, ensure realistic motion and collision behavior. Specifically, frictional forces are determined by road slope and surface properties, while gravity influences acceleration and deceleration, particularly on inclined surfaces. As the vehicle approaches the designated collision point, critical vehicle information such as speed, position, and orientation at the moment of impact are captured and the post-collision trajectory is recorded and generated. This generated post-collision trajectory data provides a robust foundation for fine-tuning the Llama model [11], significantly enhancing its ability to predict physically realistic post-collision trajectories in various scenarios.
III-B3 Fine-tuning AccidentLLM
After obtaining the collision dataset from the previous process, we fine-tune the Llama model [11] via LoRA [22] to specialize in predicting novel post-collision trajectories, resulting in a model we term AccidentLLM. This supervised fine-tuning optimizes the LLM for precise collision trajectory prediction in complex scenarios, responding to the user-provided accident descriptions.
To optimize AccidentLLM for precise trajectory prediction, we apply supervised fine-tuning using an $L_{1}$ loss function that measures the absolute discrepancy between predicted and ground truth trajectories across both spatial and rotational dimensions. The trajectory matching loss is defined as:
$L_{\text{traj}}=\frac{1}{N}\sum_{i=1}^{N}\left(\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}^{*}\|+\|\hat{\boldsymbol{\theta}}_{i}-\boldsymbol{\theta}_{i}^{*}\|\right)$ (1)
where $\hat{\mathbf{p}}_{i}=[\hat{p}_{x,i},\hat{p}_{y,i},\hat{p}_{z,i}]^{\top}$ and $\mathbf{p}_{i}^{*}$ represent the predicted and ground-truth positions at time step $i$, respectively, while $\hat{\boldsymbol{\theta}}_{i}$ and $\boldsymbol{\theta}_{i}^{*}$ denote the predicted and true rotational vectors. $N$ is the total number of time steps in the trajectory sequence. This formulation ensures that both translational and rotational components are jointly optimized, allowing AccidentLLM to capture complex post-collision dynamics and produce physically consistent trajectory predictions in diverse crash scenarios.
III-C Collision Video Generation
III-C1 Foreground Processing
As shown in Figure 2 (a), the vehicle motion information extracted from the user-provided accident description undergoes Pre-Collision Trajectory Planning, which identifies the involved vehicle models and generates both the pre-collision trajectories and key collision parameters, such as vehicle speed and collision type (e.g., rear-end collision). These results are then passed to AccidentLLM, which infers physically realistic post-collision trajectories to complete the full motion paths of the vehicles involved. Based on these complete trajectories, foreground images of the physically realistic vehicle collision are rendered.
III-C2 Background Processing
As shown in Figure 2 (b), the background processing consists of two phases: viewpoint adjustment and background rendering. In the viewpoint adjustment phase, we first parse the intended video viewpoint movements, extracted from the user-provided accident description, into positional and angular shift parameters corresponding to the target viewpoint. These shift parameters are then converted into the appropriate extrinsic transformation matrix and combined with the initial parameters to yield an updated viewpoint configuration. In the background rendering phase, if a novel viewpoint needs to be rendered, we apply the scene reconstruction method from ChatSim [50]. Otherwise, the original frame from the Waymo Open Dataset is used as the background. Specifically, ChatSim’s scene reconstruction method incorporates multi-camera alignment and luminance consistency rendering techniques, ensuring that the generated background seamlessly and naturally aligns with the vehicle’s viewpoint.
III-C3 Video Composition
In the final video composition stage, we employ alpha channel compositing techniques to finely adjust transparency between the foreground and background images, ultimately generating a visually and physically realistic collision scenario.
IV Experiments
To evaluate AccidentSim, we utilize the Waymo Open Dataset and NHTSA accident reports (Section IV-A). Our experiments focus on three core aspects: (1) Trajectory Generation Quality and Physical Consistency (Section IV-B) confirms the kinematic accuracy and semantic alignment of the generated post-collision trajectories. (2) Quality Evaluation for Generated Video (Section IV-C) demonstrates AccidentSim’s superior visual fidelity and physical realism over existing baselines. (3) Collision Reduction by AccidentSim (Section IV-D) validates the practical utility of our synthesized videos in significantly reducing collision rates for autonomous driving systems.
IV-A Datasets
We conducted experiments on the Waymo Open Dataset [44] and real-world accident reports maintained by the NHTSA. The Waymo Open Dataset contains data from 1,150 driving scenarios, which is suitable for covering the diverse scenarios required by AccidentSim. Each scenario in Waymo is recorded for 20 seconds and includes synchronized data from five LiDAR sensors and five high-resolution pinhole cameras. The NHTSA accident reports encompass a wide range of vehicle models, collision types, road types, and traffic conditions, making them suitable for use with AccidentSim.
Figure 3: Videos generated by AccidentSim in comparison with baseline methods. For a fair comparison, all methods are given the same user-provided accident description to depict the vehicle collision and driving scenario. Vehicles involved in the collision are boxed. The results show that AccidentSim follows the user description more accurately while achieving physically realistic collisions.
TABLE I: The accuracy of the generated post-collision trajectories in terms of L2 (m) for AccidentSim compared to the baseline method across different accident scenarios. The results show that AccidentSim achieves overall lower average L2 errors, demonstrating more stable and precise trajectory predictions.
Accident Scenarios Methods Trajectory Prediction Error Physical Consistency Error
L2 (m) Impulse Momentum
[…]
AutoVFX 1.21 1.90 2.42 1.84 2.89 2.91 2.77 2.65
[…]
AutoVFX 1.02 1.43 1.82 1.42 1.93 1.87 1.76 1.84
[…]
AutoVFX 1.04 2.32 2.76 2.04 3.10 3.02 2.87 2.71
[…]
AutoVFX 1.45 1.98 2.58 2.00 3.21 3.18 2.89 2.92
[…]
TABLE II: Human study results on the language-conditioned simulation.
Methods Crash Report Attribute Description
Ours Preferred (%) $\uparrow$ Score (1–5) $\uparrow$ Ours Preferred (%) $\uparrow$ Score (1–5) $\uparrow$
TrafficGen [14] 95.68 1.49 94.51 2.63
MotionCLIP [46] 96.49 1.71 95.71 2.20
LCTGen [45] 89.53 3.93 87.29 4.10
[…]
TABLE III: Quality assessment results for videos generated by AccidentSim compared to baseline methods, using VBench scores  [23], human evaluation  [5], and GPT-4o evaluation  [5]. The best value is in bold and the second-best is underlined. The results show that AccidentSim achieves state-of-the-art performance across the evaluation metrics.
Methods VBench Evaluation [23] Human Evaluation [5] GPT-4o Evaluation [5]
[…]
Kling 2.0 [27] 0.991 0.697 2.350 3.186 2.872 0.542 0.758 0.571
Gen-4 [40] 0.993 0.686 2.456 3.497 2.403 0.614 0.891 0.492
Jimeng 3.0 [3] 0.996 0.727 3.011 3.703 3.219 0.724 0.855 0.724
Cosmos [50] 0.985 0.637 2.578 3.142 2.894 0.451 0.703 0.510
AutoVFX [21] 0.990 0.701 2.983 3.286 3.450 0.521 0.768 0.682
[…]
TABLE IV: Collision rates for AccidentSim and baseline methods across diverse accident scenarios. The best value is in bold and the second-best is underlined. The results show that AccidentSim helps significantly reduce vehicle collisions by achieving lower collision rates.
Methods Dual-Vehicle Scenarios Triple-Vehicle Scenarios Multi-Vehicle Scenarios Avg.
Straight Obstacle Unprotected Left-Turn Right- Turn Lane Changing Straight Obstacle Unprotected Left-Turn Right- Turn Lane Changing Straight Obstacle Unprotected Left-Turn Right- Turn Lane Changing
Pre Fine-tuning 0.48 0.67 0.46 0.58 0.52 0.59 0.47 0.63 0.43 0.60 0.51 0.68 0.552
Learning-to-collide [9] 0.12 0.00 0.37 0.51 0.22 0.13 0.40 0.52 0.28 0.25 0.39 0.46 0.304
AdvSim [47] 0.23 0.05 0.41 0.53 0.31 0.10 0.38 0.45 0.28 0.17 0.35 0.58 0.320
Carla Scenario Generator [7] 0.22 0.22 0.19 0.39 0.17 0.26 0.17 0.48 0.12 0.29 0.10 0.48 0.258
Adversarial Trajectory Optimization [66] 0.14 0.00 0.23 0.30 0.21 0.04 0.28 0.21 0.23 0.02 0.31 0.18 0.179
ChatScene [64] 0.03 0.10 0.01 0.11 0.09 0.21 0.06 0.24 0.05 0.16 0.09 0.19 0.112
[…]
Figure 4: Vehicle collision scenarios across four typical road types. The results show that AccidentSim generates vehicle collision videos with high visual fidelity and physically realistic collision trajectories.
IV-B Trajectory Generation Quality and Physical Consistency
1) Physical Consistency and Accuracy: We evaluate the physical plausibility of AccidentLLM’s post-collision trajectories against CARLA-simulated ground truth. Trajectory Prediction Error is measured via L2 distance, and Physical Consistency Error is assessed using the MAE and RMSE of impulse and momentum changes [69]. We select AutoVFX [21], a physics-based simulation method, as our primary baseline. As shown in Table I, AccidentSim consistently outperforms AutoVFX. It achieves lower average L2 errors and significantly smaller impulse and momentum errors, demonstrating stricter adherence to physical laws like momentum conservation. Figure 3 further visualizes these physically constrained trajectories, which exhibit smoother and more realistic dynamics.
2) Text-to-Trajectory Semantic Consistency: Following the LCTGen [45] protocol, we conducted human A/B tests to measure Preference (Ours Preferred %) and Alignment (1–5 scores) on Crash Report and Attribute Description datasets. Table II demonstrates that AccidentSim significantly outperforms baselines. By fine-tuning on simulator-calibrated data, AccidentLLM effectively maps complex textual narratives to kinematically consistent trajectories.
IV-C Quality Evaluation for Generated Video
This experiment evaluates the visual fidelity and physical realism of generated vehicle collisions. Since conventional metrics (e.g., PSNR, FID, FVD) require ground-truth data, we adopt the PhysGen3D [5] protocol for reference-free evaluation. First, we use VBench [23] to assess motion smoothness and imaging quality. Second, we employ human and GPT-4o [36] evaluations to measure three key metrics: Physical Realism (PhysReal), Photorealism (PhotoReal), and Semantic Consistency (Align). Informed consent was obtained from all participants involved in the human evaluation.
Specifically, PhysReal measures adherence to physical laws, capturing dynamics like gravity and momentum; PhotoReal assesses overall visual fidelity, focusing on the absence of artifacts and the accurate depiction of lighting and shadows; and Align evaluates how accurately the generated video reflects the semantics of the user-provided accident description.
We compare AccidentSim against five state-of-the-art video generation models: Kling 2.0 [27], Gen-4 [40], Jimeng 3.0 [3], Cosmos [1], and AutoVFX [21]. For a fair comparison, all models receive identical accident descriptions and background scenes. As shown in Table III, AccidentSim achieves state-of-the-art performance in motion smoothness, physical realism, and semantic consistency, while maintaining highly competitive imaging quality. Qualitative comparisons in Figure 3 and diverse scenario results in Figure 4 further demonstrate AccidentSim’s superiority in generating physically realistic and controllable vehicle collisions.
IV-D Collision Reduction by AccidentSim
Collision reduction is the primary objective of using AccidentSim to train autonomous driving systems. This experiment evaluates whether the collision rate (CR) can be reduced by augmenting training with the collision videos generated by AccidentSim. A lower collision rate indicates that the generated videos more effectively improve these systems’ ability to avoid collisions.
Table IV reports collision rates for AccidentSim across diverse accident scenarios. For baseline comparisons, on the one hand, we compare AccidentSim with several mainstream adversarial and knowledge-driven methods for generating vehicle collision scenarios in simulated environments, including Learning-to-Collide [9], AdvSim [47], the CARLA Scenario Generator [7], Adversarial Trajectory Optimization [66], and ChatScene [64]. On the other hand, we further evaluate the effect of incorporating AccidentSim’s generated collision videos on reducing collision rates. To do this, we follow the setup from ChatScene [64] and fine-tune a model trained with Soft Actor-Critic from SafeBench [55], referred to as “Pre Fine-tuning” in Table IV, without AccidentSim’s generated collision videos for training.
V Conclusion
In this paper, we introduce AccidentSim, an innovative framework designed to generate vehicle collision videos with physically realistic collision trajectories. By extracting and utilizing physical cues and contextual information from real-world accident reports, AccidentSim effectively mitigates the critical scarcity of diverse collision videos essential for training robust autonomous driving models. Using a reliable physical simulation engine, AccidentSim ensures realistic vehicle motion in collision video generation. Leveraging a language model-driven approach, AccidentSim excels in predicting and planning novel, physically consistent collision trajectories. Experimental results demonstrate that AccidentSim generates vehicle collision videos with high visual fidelity and physical consistency. Future work will integrate vehicle deformation and additional physical cues to improve realism.